OpenStudy (anonymous):

Can somone please help me? Suppose that y = 37,000(1.005)x represents the median household income in the U.S. , where x represent the number of years after 1980. What was the median household income in 1980 (x = 0)? $37,000 $37,185 $74,000 $35,238

OpenStudy (anonymous):

since x=0, what do you know of \(\large (any number)^0= \)????

OpenStudy (anonymous):

i suck at math too, but i know this one...

OpenStudy (anonymous):

no idk anything about that lol

OpenStudy (anonymous):

do you at least have a calculator ?

OpenStudy (anonymous):

give me these values: \(\large 2^0, 5^0, 10.98^0, \pi^0, (\frac{2}{7})^0 \)